Find the expected value for the bettor of a $20 bet of one play on red in roulette. Round to the fourth as needed.
Answer #1

probability of red on roulette.P(X=20) =18/38 (snce on 18 places red is there)

P(losing) =P((X=-20)=1-18/38 =20/38

therefore expected value =20*(18/38)-20*(20/38)= -1.0526